# MnBr2 on the graphene on Ir(110) substrate: growth, structure, and super-moiré

### v1

@@ IV Methods
Ir(110) is cleaned and prepared in its unreconstructed state by cycles of 4.5 keV Xe+ ion sputtering, flash annealing to 1510 K, and subsequent cooling to 400 K in $1\times 10^{-7}$ mbar oxygen pressure.
Single-crystal Gr was grown by heating cleaned unreconstructed Ir(110) to 1510 K and exposing it to $3\times 10^{-7}$ mbar ethylene for 240 s [56]. The quality of the as-grown Gr/Ir(110) is checked by LEED and STM.
@@
LEED patterns were acquired at 300 K using electron energies ranging from 90 to 150 eV. For the MCP-LEED measurements, distortions in the reciprocal space due to the flat microchannel plate geometry were corrected using a Python script.
STM was performed at 300 K in a variable temperature and at 77 K and 1.7 K in a bath cryostat STM system. Constant-current topographies were recorded with sample bias $V_{\mathrm{b}}$ and tunneling current $I_{\mathrm{t}}$, as detailed in the respective figure captions. The STM images were analyzed and processed (plane subtraction, contrast correction) using the WSxM software [63]. STS spectra were obtained at 1.7 K with stabilization bias $V_{st}$ and stabilization current $I_{st}$ using the standard lock-in technique with modulation frequency $f_{mod}$ and modulation voltage $V_{mod}$, as specified in the captions.
Our ab initio density functional theory (DFT) [64, 65] calculations were carried out using the projector augmented wave (PAW) method [66] as implemented in the Vienna Ab initio Simulation Package (VASP) [67, 68, 69]. To account for the van der Waals interactions in the Gr/Ir(110) system, we employed the non-local vdW-DF2 correlation energy functional [70] combined with a re-optimized B86b exchange functional [71, 72]. Gr on Ir(110) was modeled using a slab comprising three Ir layers and a vacuum region of $\approx 21$ Å amounting to 350 C and 264 Ir atoms. The ground-state geometry and electronic structure were determined using a plane-wave kinetic energy cutoff of 500 eV and a threshold for Hellmann-Feynman forces of $\approx 0.005$ eV/Å. Structural relaxation was performed using $\Gamma$-point sampling of the Brillouin zone, whereas the density of states was obtained using a $\times2$ $k$-point mesh.
@@
Funding from the Deutsche Forschungsgemeinschaft (DFG) through CRC 1238 (project number 277146847, subprojects A01, C01, and B06) is acknowledged. J.F. acknowledges financial support from the DFG through project FI 2624/1-1 (project no. 462692705) within SPP 2137. Financial support by the DFG priority program SPP2244 is acknowledged W.J. (project no. 535290457). The authors acknowledge the computing time granted by the JARA Vergabegremium and provided on the JARA Partition part of the supercomputer JURECA at Forschungszentrum Jülich.

### v2

@@ IV Methods
Ir(110) is cleaned and prepared in its unreconstructed state by cycles of 4.5 keV Xe+ ion sputtering, flash annealing to 1510 K, and subsequent cooling to 400 K in $1\times 10^{-7}$ mbar oxygen pressure. Ir(111) is cleaned by cycles of 1 keV Ar+ sputtering and subsequent flash annealing to 1510 K.
Single-crystal Gr was grown by heating cleaned unreconstructed Ir(110) to 1510 K and exposing it to $3\times 10^{-7}$ mbar ethylene for 240 s [Kraus22]. On Ir(111), Gr was grown by exposing the clean Ir(111) to $1\times 10^{-7}$ mbar ethylene at 300 K for 120 s, flash annealing to 1470 K without ethylene, and again exposure to $3\times 10^{-7}$ mbar ethylene for 600 s at 1370 K [vanGastel09]. The quality of the as-grown Gr/Ir(110) or Gr/Ir(111) is checked by LEED and STM.
@@
LEED patterns were acquired at 300 K using electron energies ranging from 90 to 150 eV. For MCP-LEED measurements, distortions in the reciprocal space due to the flat microchannel plate geometry were corrected using a Python script.
STM was performed at 300 K in a variable temperature and at 77 K and 1.7 K in a bath cryostat STM system. Constant-current topographies were recorded with sample bias $V_{\mathrm{b}}$ and tunneling current $I_{\mathrm{t}}$, as detailed in the respective figure captions. The STM images were analyzed and processed (plane subtraction, contrast correction) using WSxM software [Horcas2007]. STS spectra were obtained at 1.7 K with stabilization bias $V_{st}$ and stabilization current $I_{st}$ using the standard lock-in technique with modulation frequency $f_{mod}$ and modulation voltage $V_{mod}$, as specified in the captions.
Our ab initio density functional theory (DFT) [PR136_B864, PR140_A1133] calculations were carried out using the projector augmented wave (PAW) method [PRB50_17953] as implemented in the Vienna Ab initio Simulation Package (VASP) [PRB47_558, PRB54_11169, PRB59_1758]. To account for the van der Waals interactions in the Gr/Ir(110) and MnBr2/Gr/Ir(110) systems, we employed the non-local vdW-DF2 correlation energy functional [PRB82_081101R] combined with a re-optimized B86b exchange functional [PRB89_121103R, JCP86_7184]. Gr on Ir(110) was modeled using a slab comprising three Ir layers and an in-plane unit cell of $\approx 32.86\times 30.12$ Å amounting to 614 atoms (350 C and 264 Ir). The MnBr2/Gr/Ir(110) system was simulated by using an in-plane unit cell of $\approx 7.04\times 28.46$ Å that contains 192 atoms (i.e., 14 Mn, 28 Br, 84 C and 66 Ir). The ground-state geometry and electronic structure were determined using a plane-wave kinetic energy cutoff of 500 eV and a threshold for Hellmann-Feynman forces of $\approx 0.005$ eV/Å.
@@
Funding from the Deutsche Forschungsgemeinschaft (DFG) through CRC 1238 (project number 277146847, subprojects A01, C01, and B06) is acknowledged. J.F. acknowledges financial support from the DFG through project FI 2624/1-1 (project no. 462692705) within SPP 2137. W.J. acknowledges financial support by the DFG priority program SPP2244 (project no. 535290457). The authors acknowledge the computing time granted by the JARA Vergabegremium and provided on the JARA Partition part of the supercomputer JURECA at Forschungszentrum Jülich.
@@
VI Supplementary Information
Supplementary Note 1. Apparent height of MnBr2
Figure 1: STM topograph of MnBr2 on Gr/Ir(110). Lower panel shows the height profile along the black line in the STM topograph. STM image is obtained at 300 K with (b) $V_{\mathrm{b}}$ = -2 V, 50 pA. Image size: 200 nm $\times$ 160 nm
Supplementary Note 2. Comparison of single layer MnBr2 grown on Gr/Ir(111) and Gr/Ir(110)
Figure 2: STM topograph of MnBr2 grown at 400 K on (a) Gr/Ir(111) and (b) Gr(Ir(110). (b) is identical to Figure 1(b) of the main manuscript and is reproduced here for direct comparison. Tunneling parameters are for (a) $V_{\mathrm{b}}$ = 3.5 V, 20 pA. and for (b) $V_{\mathrm{b}}$ = -2 V, 50 pA. Image size for (a) and (b) is 180 nm $\times$ 200 nm.
Supplementary Note 3. Comparison of Gr/Ir(110) with MnBr2/Gr/Ir(110)
Figure 3: (a) Moiré resolution STM topography of Gr/Ir(110) and (b) its corresponding FFT. (c) Moiré resolution STM topography of MnBr2/Gr/Ir(110) and (c) its corresponding FFT. The Fourier components encircled blue in (b) and (d) are identical. (c) and (d) are Figures 5(a) and (b) of the paper and are reproduced for direct comparison. STM imaging parameters are (a) $V_{\mathrm{b}}$ = -1 V, 100 pA and (c) $V_{\mathrm{b}}$ = -1 V, 50 pA. STM image size is 40 nm $\times$ 40 nm.
Supplementary Note 4. Contour of constant total charge density above Gr/Ir(110)
Figure 4: (a) Ball model of atomic arrangement in the supercell. Green gold balls: bottom-layer Ir atoms; white balls: top-layer Ir atoms; dark gray balls: C atoms. A pink rhombus indicates the supercell size in the DFT calculations. (b) The total charge density plotted above the surface for an isosurface value of $2.1\times 10^{-10}$ electrons/Å3. The grayscale encodes the variation in isosurface height. The periodicity of the Ir(110) top layer substrate rows with their 3.84 Å spacing is prominent.
Supplementary Note 5. DFT simulation of MnBr2/Gr/Ir(110)
Figure 5: (a) Top view of the relaxed structure of MnBr2 on Gr/Ir(110). The size of supercell is indicated by black rectangle. It is smaller by a factor of $\approx 4$ compared to the supercell used to model the Gr/Ir(110) moiré in Figure 6(d) of the main paper. (b),(c) Side views along (b) $[001]$ direction and (c) $1\bar{1}0]$ direction. Black lines indicate boundaries of supercell. (d) The charge density difference is plotted above the surface for an isosurface value $4.9\times 10^{-8}$ e/Å3. Yellow indicates charge density accumulation and blue charge density depletion.
DFT calculations for the entire super-moiré unit cell are impossible. Even including the MnBr2 layer in the unit cell as shown in Figure 9(a) is computationally very expensive and practically not feasible. However, to get an insight into the interaction that occurs between MnBr2 and the Gr/Ir(110) substrate we can use a smaller supercell that correctly captures the chemistry at the interface of MnBr2 and Gr/Ir(110). The supercell used in our calculations is indicated by a black rectangle in Figure S5(a) and contains 14 Mn, 28 Br, 84 C, and 66 Ir atoms. The relative lattice parameters with respect to their nominal values are 0.92 for Ir, 0.96 for Gr, and 1.05 for Mn. We note here that the choice of the unit cell implies a compressive stress for Gr/Ir(110) substrate and tensile stress for the MnBr2 layer.
[…]